OpenStudy (anonymous):

Find the distance between these lines: y = -2x + 6 and y = -2x - 4

OpenStudy (anonymous):

I know I have to find the slope of what is perpendicular to the lines, which is 1/2, but I don't know what to do next.

OpenStudy (earthcitizen):

equate y and find x

OpenStudy (anonymous):

Put it in standard form? y=-2x+6 2x + y = 6 2x = 6 -y x=3-1/3y

OpenStudy (anonymous):

How's that supposed to help?

OpenStudy (anonymous):

How do I do this? -2x + 6 = -2x - 4 10 = 0?

OpenStudy (earthcitizen):

use x value to find y co-ordinates

OpenStudy (accessdenied):

this would be the length of a perpendicular line drawn between them... we could find an arbitrary equation for the perpendicular line (point slope, any y-intercept), find where it intersects each line (finding where our new line crosses each line), and then distance formula to find the specific distance.

OpenStudy (accessdenied):

not point slope, slope-intercept -- i'm mixing myself up. :P

OpenStudy (anonymous):

Still not catching on... how exactly would I do this?

OpenStudy (accessdenied):

do you remember how to do systems of equations? basically, you'd be finding a perpendicular line to those lines (they're parallel lines), finding the solution for both pairs of perpendicular lines, and then using the points you found in distance formula. This would be the most intuitive way of doing it, to me something like this: (not using your equations, just some random lines)